# Personalized Enhanced Federated Multi-View Clustering via Heat-Kernel Tensor Decomposition

### v2

@@ 3 Mathematical Preliminaries
This section presents a thorough mathematical formulation of the fundamental concepts and notations employed in this study. The section commences with a review of the principles of federated learning and clustering. This is followed by an examination of multi-view clustering techniques. Subsequently, the section explores heat kernel methods and their applications in quantum-inspired learning. Finally, the discussion will turn to tensor decomposition methods relevant to the proposed framework.
@@
3.1 Federated Learning and Clustering
Federated learning, pioneered by McMahan et al. [mcmahan2017communication], enables collaborative machine learning across distributed clients without centralizing raw data. The seminal FedAvg algorithm established the foundation for aggregating local model updates through weighted averaging. However, traditional federated learning primarily focuses on supervised tasks, with federated clustering receiving comparatively less attention despite its practical importance in distributed data analytics.
Konečnỳ et al. [konevcny2016federated] introduced federated optimization as a method to train a centralized model using decentralized data stored across numerous devices, emphasizing communication efficiency due to limited connectivity and privacy concerns. They proposed modifications to existing algorithms to adapt them to federated optimization settings. Li et al. [li2020acceleration] adopted the foundational ideas of federated optimization and extended them by integrating advanced techniques like gradient compression, acceleration, and variance reduction to address communication bottlenecks and improve the scalability of federated learning systems
Early work on federated clustering by Dennis et al. [dennis2021heterogeneity], Ghosh et al. [ghosh2002consensus], and Makhija et al. [makhija2022federated] demonstrated the feasibility of distributed k-means and hierarchical clustering. These approaches typically involve local clustering followed by global aggregation of cluster centers or assignments. However, these methods assume homogeneous data distributions and fail to address the inherent heterogeneity in federated environments.
Recent advances have introduced personalized federated clustering to address client heterogeneity. Caldarola et al. [caldarola2021cluster] proposed adaptive weight schemes for heterogeneous clustering, while Li et al. [li2022hpfl] developed meta-learning approaches for personalized cluster initialization. These works, however, are limited to single-view data and do not leverage the rich multi-modal information available in modern federated systems.
Privacy-preserving aspects of federated clustering have been addressed through differential privacy mechanisms [kairouz2021advances] and secure aggregation protocols [bonawitz2017practical]. Geyer et al. [geyer2017differentially] demonstrated how differential privacy can be integrated into federated clustering while maintaining acceptable utility. However, these privacy mechanisms often come at the cost of clustering accuracy, particularly in high-dimensional settings.
The communication efficiency challenge in federated clustering has been tackled through gradient compression [konecny2016federated], sparsification techniques [aji2017sparse], and model quantization [reisizadeh2020fedpaq]. While these methods reduce communication overhead, they typically do not exploit the inherent structure of multi-view tensor data for more efficient representation and transmission.
3.2 Multi-View Clustering
Multi-view clustering leverages complementary information from multiple data representations to improve clustering performance. Early work by Bickel and Scheffer [bickel2004multi] established the theoretical foundations for multi-view learning, demonstrating superior performance when views provide complementary information about the underlying data structure.
Classical multi-view clustering approaches can be categorized into three main paradigms: co-training based methods, subspace learning approaches, and consensus clustering techniques. Co-training methods, such as the work by Kumar et al. [kumar2011co], iteratively refine clustering results by leveraging inter-view consistency. Subspace learning approaches, including Canonical Correlation Analysis (CCA) based methods [hotelling1992relations] and shared subspace clustering [wang2019multiple], project multiple views into a common representation space for unified clustering.
Graph-based multi-view clustering has gained significant attention due to its ability to capture complex inter-view relationships. Wang et al. [wang2019gmc] proposed multi-view graph clustering with adaptive neighbor learning, while Nie et al. [nie2016parameter] developed parameter-free auto-weighted multiple graph learning. These methods construct view-specific graphs and learn optimal view weights for clustering.
Deep learning approaches have revolutionized multi-view clustering through end-to-end representation learning. Zhang et al. [zhang2019ae2] introduced auto-encoder based multi-view clustering, while Zhu et al. [zhu2019multi] developed deep multi-view clustering networks with view correlation discovery. However, these methods typically require centralized data access and are not designed for federated environments.
Fuzzy clustering extensions to multi-view scenarios have been explored by Jiang et al. [jiang2014collaborative] and Wang et al. [wang2017multi], introducing weighted aggregation schemes and view-specific fuzzification parameters. These soft clustering approaches provide uncertainty quantification but lack the scalability and privacy considerations required for federated deployment.
The mathematical foundations of multi-view clustering build upon single-view clustering formulations. In this study, we focus on soft clustering approaches that provide probabilistic cluster assignments, enabling better uncertainty quantification in federated environments. The objective function of standard fuzzy c-means (FCM) or soft clustering can be expressed as:
@@ 3.2 Multi-View Clustering
As displayed in Eq. 2, the memberships of the $h$-th view data in $c$ clusters $\mu_{ik}^{h}$ is computed locally. Usually, to get a global memberships of data view $s$, we have an additional step called as aggregation of collected memberships across data views $s$. There are multiple ways to aggregate the memberships of $s$ data views, i.e., using the averagation approach or taking a product of memberships of the $i$-th data point across all views $s$ in clusters $c$ and divide it with the number of data points. If we consider a weight factor $v_{h}$, we can multiply each weight factors with the memberships and divide it by the number of data views, etc. As we focused on a client’ multi-view data, we simplify the computation for each view memberships by directly implementing the global memberships from an early iteration. In this way, we can modify Eq. 2 in the following way
@@
where $\delta_{ij}^{h}$ is a heat kernel coefficient (HKC) to control the Euclidean distance between data points $x_{ij}^{h}$ and cluster centers $a_{kj}^{h}$ across data views in kernel space. In this way, we can modify Eq. 3 in the following way
$J_{KMVC}(U^{*},A^{h})=\sum_{h=1}^{s}\sum_{i=1}^{n}\sum_{k=1}^{c}(\mu_{ik}^{*})^{m}\text{KED}(x_{ij}^{h},a_{kj}^{h})$ (5)
@@
As illustrated in Eq. 5, it is imperative to prioritize all data views equally. This approach is not conducive to the efficient resolution of multiple resources data. Given that divergent perspectives correspond to disparate manifestations of feature behavior, it becomes imperative to allocate a distinct weight, denoted as $v_{h}$, to each data view, denoted as $x_{i}^{h}$. It is imperative to underscore that this allocation must be completed, in accordance with the principle that $\sum_{h=1}^{s}v_{h}=1$. By incorporating the weight factors, denoted by $v_{h}$, into the analysis, a modification to the Eq. 5 can be formulated as follows:
@@
where $\alpha$ is an exponent parameter to control the distribution of data views during the clustering processes.
3.3 Heat Kernel Methods and Quantum-Inspired Learning
Heat kernel methods have emerged as powerful tools for capturing local geometric structure in high-dimensional data spaces. Originally developed in differential geometry and quantum field theory, heat kernels provide a natural way to encode multi-scale geometric information through diffusion processes on manifolds.
The heat kernel, defined as the fundamental solution to the heat equation, has found extensive applications in machine learning through kernel methods and manifold learning. Ham et al. [ham2004kernel] introduced diffusion maps that leverage heat kernel properties for dimensionality reduction, while Coifman and Lafon [coifman2006diffusion] established theoretical foundations for diffusion-based data analysis.
In clustering applications, heat kernel methods have been used to capture local neighborhood structure and improve cluster boundary detection. Zelnik-Manor and Perona [zelnik2004self] developed self-tuning spectral clustering using local heat kernel scales, while von Luxburg [von2007tutorial] provided comprehensive analysis of heat kernel properties in graph-based clustering.
Quantum-inspired machine learning has gained attention for its ability to leverage quantum mechanical principles in classical algorithms. Quantum kernels, introduced by Havlíček et al. [havlivcek2019supervised], enable the computation of quantum feature maps for enhanced pattern recognition. The connection between quantum field theory and machine learning has been explored through path integral formulations [carleo2017solving] and quantum neural networks [biamonte2017quantum].
The integration of heat kernel coefficients with clustering algorithms represents a natural extension of kernel methods to soft clustering scenarios. By incorporating heat diffusion properties, clustering algorithms can better capture local data manifold structure and improve robustness to noise and outliers.
3.4 Federated Multi-View Learning and Privacy Preservation
The intersection of federated learning and multi-view clustering presents unique challenges that are not adequately addressed by existing literature. Most federated learning approaches assume single-view data and homogeneous client populations, while multi-view clustering methods typically require centralized data access.
Recent work on federated multi-view learning includes approaches by Chen et al. [chen2022federated] for federated multi-view representation learning and Liu et al. [liu2023federated] for privacy-preserving multi-view clustering. However, these methods do not leverage tensor decomposition for efficient representation and communication.
Privacy preservation in federated multi-view settings requires careful consideration of information leakage across views and clients. Differential privacy mechanisms must be adapted to handle the multi-view tensor structure while maintaining clustering utility. Secure aggregation protocols need to be extended to handle tensor factor aggregation efficiently.
The communication efficiency challenge is particularly acute in federated multi-view scenarios due to the high-dimensional nature of multi-view tensor data. Existing compression techniques are not optimized for tensor structures and may destroy important multilinear relationships. This motivates the need for tensor-aware compression and aggregation schemes.
The subsequent discourse will encompass the tensorization and activation of equation Eq. 6. In the following sections, the open challenges will be addressed by means of a proposed heat-kernel-enhanced tensorized framework in federated environments.
@@ 3.5 Tensor Decomposition Methods
Tensor decomposition has emerged as a powerful paradigm for analyzing high-dimensional multi-modal data, providing compact representations that capture multilinear relationships inherent in complex datasets. The field has evolved from classical matrix factorization techniques to sophisticated tensor algebra frameworks capable of handling arbitrary N-way data structures.
Canonical Polyadic Decomposition (CPD), also known as CANDECOMP/PARAFAC, represents the foundational tensor decomposition method. Introduced independently by Carroll and Chang [carroll1970analysis] and Harshman [harshman1970foundations], CPD decomposes a tensor into a sum of rank-one tensors, providing a unique factorization under mild conditions. The uniqueness property of CPD makes it particularly valuable for component analysis and factor identification in multi-way data.
Tucker decomposition, proposed by Tucker [tucker1966some], provides a more flexible alternative through its core tensor and factor matrices structure. Unlike CPD, Tucker decomposition allows different ranks along each mode, making it suitable for data with varying complexity across dimensions. The Higher-Order SVD (HOSVD) algorithm by De Lathauwer et al. [de2000multilinear] established efficient computational procedures for Tucker decomposition, enabling its application to large-scale problems.
Recent advances in tensor decomposition have focused on addressing practical challenges such as missing data, noise robustness, and computational scalability. Tensor completion methods, including work by Liu et al. [liu2012tensor] and Yuan and Zhang [yuan2016tensor], enable handling of incomplete tensors through low-rank assumptions. Robust tensor decomposition techniques [huang2014robust] provide resilience against outliers and corrupted entries.
Non-negative tensor factorization (NTF) has gained attention for applications requiring interpretable, parts-based representations. Lee and Seung’s pioneering work [lee2000algorithms] on non-negative matrix factorization was extended to tensors by Shashua and Hazan [shashua2005non], enabling decompositions with non-negativity constraints that often yield more interpretable factors.
The integration of tensor decomposition with machine learning has led to tensorized neural networks [novikov2015tensorizing], tensor-based kernel methods [signoretto2011kernel], and multi-task learning frameworks [wimalawarne2014multitask]. These approaches leverage tensor structure to reduce model complexity while maintaining representational power.
In the context of clustering, tensor-based methods have shown particular promise for multi-view and multi-way data analysis. Zhou et al. [cichocki2015tensor] developed tensor-based multiway arrays for bid data sets, while Benjamin et al. [benjamin2024tensor] proposed tensorized possibilistic c-means for high-dimensional clustering. However, these methods are not designed for federated environments and lack the privacy-preserving mechanisms required for distributed deployment.
A tensor is a higher-order array that represents multi-dimensional data structures. Scalar is a zero-order tensor; Vector is a 1-dimensional array, also known as tensor of order one and denoted by boldface lowercase letter, e.g., $\mathbf{x}$; A matrix $\mathbf{X}$ is a 2-way array or tensor of order two with two indices $(m,n)$, where $m$ indicates the row dimension and $n$ indicates the column dimension; While a tensor $\mathcal{X}$ is a generalization of matrices with three or more indices $(i,j,k,\ldots)$, considered as 3-way arrays or higher-way or N-way arrays (higher-order) tensors.
When $X\in\mathbb{R}^{n\times d_{h}\times s}$ represents multi-view data in tensor space, the notation $\mathcal{X}_{ijh}$ denotes the score or element of the $i$-th data point on the $j$-th feature component at the $h$-th view. Figure 2 illustrates the representation of multi-view data in tensor space. As clearly shown, multi-view data can be viewed as a 3-dimensional tensor $\mathcal{X}=\{x_{ijh}\}_{i=1,j=1,h=1}^{n,d_{h},s}$ or equivalently as a set of 2D matrices $\mathcal{X}=\{X^{1},\ldots,X^{s}\}$ with size $n\times d_{h}$.
Figure 2: The Illustration of MV Data in Tensor Space
Tensor Operations and Decomposition Preliminaries: The elements of $N$-way arrays can be reordered into matrices or vectors through matricization and vectorization operations. The matricization is reordering tensor’ elements into a matrix, while vectorization is reordering tensor’ elements into a vector. Canonical polyadic decomposition (CPD) or parallel factors (PARAFAC) [harshman1994parafac, carroll1970analysis], Tucker decomposition (TD), and non-negative tensor factorization (NTF) are well-known approaches to reorder a tensor into a matrix or vector. PARAFAC or CPD holds multilinearity assumptions such as handle a sum of rank-one tensors by targeting the same lengths. In this sense, the original PARAFAC decomposition can reveal uncorrelated (orthogonal) and correlated (oblique) factors in system-variation data matrices, but failed to describe correlated factors when applied to the cross-product matrices computed from one data [harshman1972parafac2, kiers1999parafac2]. PARAFAC2, on the other hand has the advantageous to handle irregular tensors by relaxing the canonical polyadic (CP) model [harshman1972parafac2]. PARAFAC2 enables the three-mode model, often implemented into subspace clustering algorithms taking advantageous of correlations between pair of factors. The PARAFAC2 normalize the generalization of PARAFAC to deal both with uncorrelated and correlated factors.
@@ 3.5.1 The Basic Background of Tensors
Figure 3: The Illustration of Rank-One Third-Order Tensor $\mathcal{A}=x\circ y\circ z$
@@
where $i_{n}=j_{m}=q_{p}$ and the tensor $\mathcal{C}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n-1}\times G_{n+1}\times\ldots\times G_{N}\times J_{1}\times J_{2}\times\ldots\times J_{m-1}\times J_{m+1}\times\ldots\times J_{M}}$.
@@ 3.5.2 The Higher Order Singular Value Decomposition (HOSVD)
Tensor factorization is a method for reducing a large volume of data to a smaller set of data points. This is achieved by leveraging the collinearity of its heterogeneous data [wang2017tensor]. CANDECOMP/PARAFAC is a technique for decomposing data into its constituent parts. It is one of the most widely used tensor decomposition techniques, as referenced in the works of Kolda et al. (2009) [kolda2009tensor]. and Kiers (2000) [kiers2000towards]. Conversely, the HOSVD represents an extension of the matrix singular value decomposition (SVD). In order to provide a more thorough explanation of the higher-order SVD (HOSVD), this paper will draw upon certain concepts from Wang and Song’s work on two-mode tensor orders $N$ [wang2017tensor].

### v3

@@ 2 Mathematical Preliminaries
This section presents a thorough mathematical formulation of the fundamental concepts and notations employed in this study. This includes an mathematical overview of fuzzy c-mean clustering, multi-view clustering, kernel Euclidean distance, tensor algebra, and tensor decomposition techniques, specifically PARAFAC2 and Tucker decomposition. These mathematical preliminaries provide the necessary foundation for understanding the proposed federated multi-view clustering framework.
2.1 Multi-View Clustering
@@
Soft Clustering: The mathematical foundations of multi-view clustering build upon single-view clustering formulations. In this study, we focus on soft clustering approaches that provide probabilistic cluster assignments, enabling better uncertainty quantification in federated environments. The objective function of standard fuzzy c-means (FCM) or soft clustering can be expressed as:
@@
As displayed in Eq. 2, the memberships of the $h$-th view data in $c$ clusters $\mu_{ik}^{h}$ is computed locally. Usually, to get a global memberships of data view $s$, we have an additional step called as aggregation of collected memberships across data views $s$. There are multiple ways to aggregate the memberships of $s$ data views, i.e., using the averagation approach or taking a product of memberships of the $i$-th data point across all views $s$ in clusters $c$ and divide it with the number of data points. If we consider a weight factor $v_{h}$, we can multiply each weight factors with the memberships and divide it by the number of data views, etc. As we focused on a client’ multi-view data, we simplify the computation for each view memberships by directly implementing the global memberships $U^{*}$ from an early iteration. In this way, we can modify Eq. 2 in the following way
@@
where $\delta_{ij}^{h}$ is a heat kernel coefficient (HKC) to control the Euclidean distance between data points $x_{ij}^{h}$ and cluster centers $a_{kj}^{h}$ across data views in kernel space. By substituting the Euclidean distance in Eq. 3 with the KED defined in Eq. 4, we obtain the following objective function for alternative multi-view clustering (AMVC):
$J_{AMVC}(U^{*},A^{h})=\sum_{h=1}^{s}\sum_{i=1}^{n}\sum_{k=1}^{c}(\mu_{ik}^{*})^{m}\text{KED}(x_{ij}^{h},a_{kj}^{h})$ (5)
@@
As illustrated in Eq. 5, it is imperative to prioritize all data views equally. This approach is not conducive to the efficient resolution of multiple resources data. Given that divergent perspectives correspond to disparate manifestations of feature behavior, it becomes imperative to allocate a distinct weight, denoted as $v_{h}$, to each data view, denoted as $x_{i}^{h}$. It is imperative to underscore that this allocation must be completed, in accordance with the principle that $\sum_{h=1}^{s}v_{h}=1$. By incorporating the weight factors, denoted by $v_{h}$, into the analysis, a modification to the Eq. 5 is proposed, resulting in the following objective function for the efficient kernel multi-view clustering (E-KMVC):
@@
where $\alpha$ is an exponent parameter to control the distribution of data views during the clustering processes. Eq. 6 serves as the foundational objective function for the proposed efficient federated kernel multi-view clustering (E-FKMVC) framework. This framework is designed to address the challenges associated with multi-view data clustering in federated learning environments, particularly focusing on data heterogeneity, privacy preservation, and communication efficiency.
@@ 2.2 Tensor Decomposition Methods
Tensor Notations: In this subsection, we introduce some basic notations and definitions of tensor algebra. Most of the overviews and notations are borrowed from Kolda et al. (2009) [kolda2009tensor].
A tensor is a multi-dimensional array that generalizes the concepts of scalars (0th-order tensors), vectors (1st-order tensors), and matrices (2nd-order tensors) to higher dimensions. An $N$-way or $N$-th order tensor is denoted by a calligraphic letter, e.g., $\mathcal{X}\in\mathbb{R}^{I_{1}\times I_{2}\times\ldots\times I_{N}}$, where $I_{n}$ represents the size of the tensor along the $n$-th mode. The elements of the tensor are accessed using multiple indices, e.g., $\mathcal{X}_{i_{1},i_{2},\ldots,i_{N}}$. Consider a multi-view data $\mathcal{X}=\{x_{ijh}\}_{i=1,j=1,h=1}^{n,d_{h},s}$ or equivalently as a set of 2D matrices $\mathcal{X}=\{X^{1},\ldots,X^{s}\}$ with size $n\times d_{h}$, where $i$ indicates the data point index, $j$ indicates the feature index, and $h$ indicates the view index. Each slice along the third mode represents data from a specific view (see Fig. 3).
Figure 3: Tensor Representation of Multi-View Data with Different Modalities. The number of data points $n$ is represented along the first mode, the feature dimensionality $d_{h}$ of each view is represented along the second mode, and the number of views $s$ is represented along the third mode. Each slice along the third dimension corresponds to a different view of the same data points, capturing complementary information across modalities.
@@ 2.2.1 The Basic Background of Tensors
Assume that $\mathcal{A}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{n}\times\ldots\times G_{N}}$ is an $N$-th order tensor, where $G_{n}$ represents the size of the tensor along the $n$-th mode. The elements of the tensor are accessed using multiple indices, e.g., $\mathcal{A}_{g_{1},g_{2},\ldots,g_{n}}$.
@@
Figure 4: The Illustration of Rank-One Third-Order Tensor $\mathcal{A}=x\circ y\circ z$. The $(i,j,k)$ element of $\mathcal{A}$ is given by $a_{ijk}=x_{i}y_{j}z_{k}$. In terms of element wise, the 3-way tensor can be written as $a_{ijk}\approx\sum_{r=1}^{R}x_{ir}y_{jr}z_{kr}$.
@@
where $i_{n}=j_{m}=q_{p}$ and the tensor $\mathcal{C}\in\mathbb{R}^{G_{1}\times\ldots\times G_{n-1}\times G_{n+1}\times\ldots\times G_{N}\times J_{1}\times J_{2}\times\ldots\times J_{m-1}\times J_{m+1}\times\ldots\times J_{M}}$.
@@ 2.2.2 The Higher Order Singular Value Decomposition (HOSVD)
Tensor factorization is a method for reducing a large volume of data to a smaller set of data points. This is achieved by leveraging the collinearity of its heterogeneous data [wang2017tensor]. CANDECOMP/PARAFAC is a technique for decomposing data into its constituent parts. It is one of the most widely used tensor decomposition techniques, as referenced in the works of Kolda et al. (2009) [kolda2009tensor]. and Kiers (2000) [kiers2000towards]. Tucker decomposition is another popular tensor decomposition technique that decomposes a tensor into a core tensor and factor matrices. It was first introduced by Tucker (1966) [tucker1966some] and has been widely used in various applications. Higher-Order Singular Value Decomposition (HOSVD) is a generalization of the matrix singular value decomposition (SVD) to higher-order tensors. It was introduced by De Lathauwer et al. (2000) [de2000multilinear] and has been widely used in various applications, including image processing, signal processing, and machine learning. In order to provide a comprehensive understanding of tensor decomposition techniques, we present the following three widely used methods:
The PARAFAC decomposition [carroll1970analysis] of a tensor $\mathcal{A}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{N}}$ is defined as
$\mathcal{A}\approx\sum_{r=1}^{R}\lambda_{r}a_{r}^{(1)}\circ a_{r}^{(2)}\circ\ldots\circ a_{r}^{(N)}$ (15)
where $\lambda_{r}$ is the weight of the $r$-th rank-one tensor, and $a_{r}^{(n)}\in\mathbb{R}^{G_{n}}$ is the factor vector along the $n$-th mode. The minimum value of $R$ that satisfies Eq. 15 is defined as the rank of tensor $\mathcal{A}$.
The Tucker decomposition [tucker1966some] of a tensor $\mathcal{A}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{N}}$ is defined as
$\mathcal{A}\approx\mathcal{S}\times_{1}U^{(1)}\times_{2}U^{(2)}\times_{3}\ldots\times_{N}U^{(N)}$ (16)
where $\mathcal{S}\in\mathbb{R}^{R_{1}\times R_{2}\times\ldots\times R_{N}}$ is the core tensor, and $U^{(n)}\in\mathbb{R}^{G_{n}\times R_{n}}$ is the factor matrix along the $n$-th mode. The dimensions $R_{n}$ are the multilinear ranks of tensor $\mathcal{A}$.
The Higher-Order Singular Value Decomposition (HOSVD) [de2000multilinear] of a tensor $\mathcal{A}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{N}}$ is defined as
$\mathcal{A}=\mathcal{S}\times_{1}U^{(1)}\times_{2}U^{(2)}\times_{3}\ldots\times_{N}U^{(N)}$ (17)
where $\mathcal{S}\in\mathbb{R}^{G_{1}\times G_{2}\times\ldots\times G_{N}}$ is the core tensor, and $U^{(n)}\in\mathbb{R}^{G_{n}\times G_{n}}$ is the orthogonal factor matrix along the $n$-th mode. The HOSVD provides a unique decomposition of tensor $\mathcal{A}$, and the factor matrices $U^{(n)}$ are obtained by performing SVD on the mode-$n$ unfoldings of tensor $\mathcal{A}$.
Table 1 summarizes the key characteristics of the three tensor decomposition methods discussed above.
Table 1: Summary of Tensor Decomposition Methods: Characteristics, Components, and Applications
Method Decomposition Form Components Key Features Applications PARAFAC (CP) $\mathcal{A}\approx\sum_{r=1}^{R}\lambda_{r}a_{r}^{(1)}\circ a_{r}^{(2)}\circ\ldots\circ a_{r}^{(N)}$ Weights $\lambda_{r}$ Factor vectors $a_{r}^{(n)}$ Unique under mild conditions Fixed rank $R$ Interpretable factors Chemometrics Signal processing Neuroscience Tucker $\mathcal{A}\approx\mathcal{S}\times_{1}U^{(1)}\times_{2}U^{(2)}\times_{3}\ldots\times_{N}U^{(N)}$ Core tensor $\mathcal{S}$ Factor matrices $U^{(n)}$ Flexible multilinear ranks $(R_{1},R_{2},\ldots,R_{N})$ Non-unique decomposition Higher compression ratio Image compression Dimensionality reduction Data mining HOSVD $\mathcal{A}=\mathcal{S}\times_{1}U^{(1)}\times_{2}U^{(2)}\times_{3}\ldots\times_{N}U^{(N)}$ All-orthogonal core $\mathcal{S}$ Orthonormal matrices $U^{(n)}$ Unique decomposition Orthogonal factors Optimal truncation properties Multilinear PCA Subspace analysis Tensor initialization Complexity Comparison Method Time Complexity Space Complexity PARAFAC (CP) $\mathcal{O}(R\prod_{n=1}^{N}G_{n})$ per iteration $\mathcal{O}(R\sum_{n=1}^{N}G_{n})$ Tucker $\mathcal{O}(\prod_{n=1}^{N}R_{n}+\sum_{n=1}^{N}G_{n}R_{n})$ $\mathcal{O}(\prod_{n=1}^{N}R_{n}+\sum_{n=1}^{N}G_{n}R_{n})$ HOSVD $\mathcal{O}(\sum_{n=1}^{N}G_{n}^{2}\prod_{m\neq n}G_{m})$ $\mathcal{O}(\prod_{n=1}^{N}G_{n}+\sum_{n=1}^{N}G_{n}^{2})$